Guggenheim, Inc. offers a 8.1% coupon bond with annual payments. The yield to maturity is 5.85% and the maturity date is 9 years. What is the market price of a $750 face value bond?

Select one:

a. 721.36

b. 1064.60

c. 649.15

d. 958.70

e. 865.53
Answer #1

Annual coupon=750*8.1%=$60.75

Hence market price=Annual coupon*Present value of annuity factor(5.85%,9)+$750*Present value of discounting factor(5.85%,9)

=60.75*6.846317809+$750*0.599490408

=$865.53(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=60.75[1-(1.0585)^-9]/0.0585

=60.75*6.846317809

2.Present value of discounting factor=750/1.0585^9

=$750*0.599490408